OpenStudy (anonymous):

Find the square. (2x – 6)2

sam (.sam.):

(2x-6)^(2) Multiply each term in the first group by each term in the second group using the FOIL method. 4x^(2)-24x+36

OpenStudy (king):

4x^2+36-24x

OpenStudy (callisto):

(2x – 6)^2 = (2x – 6)(2x – 6) = 2x(2x – 6) -6(2x – 6) = 4x^2 - 12x -12x +36 = 4x^2 -24x +36 Or you can use the identity (\[(a \pm b)^2 = a^2 \pm 2ab +b^2\]